OpenStudy (gurkipal):

Anybody knows why we put R = z-1 when finding poles? Thanks

OpenStudy (anonymous):

dinkleburg

OpenStudy (ybarrap):

Every polynomial can be factored as a product of roots: $$ z^n+z^{n-1}+\cdots z^2+z+c=(z-z_1)(z-z_2)\cdots(z-z_{n-1})(z-z_n) $$ In this form, it is easy to see where the poles, \(p_i\\), and zeros, \(z_i\\), lay: $$ f(z)=\frac{(z-z_1)(z-z_2)\cdots(z-z_{n-1})(z-z_n)}{(z-p_1)(z-p_2)\cdots(z-p_{n-1})(z-p_n)} $$